# HRET: A Self-Evolving LLM Evaluation Toolkit for Korean

### v2

@@ 6 Conclusion
We introduce HRET, a self-evolving, modular evaluation toolkit tailored specifically for Korean large language models. HRET standardizes methodologies, improves reproducibility, and addresses linguistic nuances specific to Korean. Future work will include expanding benchmarks for domain-specific linguistic analyses, deeper integration with emerging LLM libraries, and enhancing visualization tools to facilitate detailed error diagnostics. Ultimately, HRET aims to establish a robust foundation for Korean NLP research, supporting sustained progress and collaboration within the community.
@@
While HRET provides a standardized evaluation framework for Korean LLMs, several limitations remain:
First, despite integrating major Korean benchmarks, coverage in specialized domains (legal, medical, scientific) remains limited. Second, the language penalizer may struggle with mixed-language outputs or code-switching common in technical contexts. Third, computational requirements for scaling techniques can be substantial, potentially limiting accessibility.
Current evaluation metrics, particularly string matching, may not fully capture Korean’s rich morphological structures. Additionally, as new Korean LLMs and benchmarks emerge, maintaining compatibility requires continuous updates. Finally, the toolkit offers limited tools for analyzing model errors specific to Korean linguistic phenomena.
We are addressing these limitations through ongoing development and welcome community contributions to enhance HRET’s capabilities.
Appendix A Usage Examples and Tutorials
@@
A.1 Advanced Backend Usage
@@
A.1.1 LiteLLM and OpenAI-Compatible APIs
@@
A.2 Evaluation Methods
A.2.1 String Match and Log Probability Evaluation
@@
A.2.2 Math Verification
Semantic equivalence checking for mathematical expressions:
results = evaluator.run(
model="huggingface",
model_params={"cot": True}, # Enable Chain-of-Thought
dataset="hrm8k",
evaluation_method="math_verify",
evaluator_params={"extract_final_answer": True}
A.3 Advanced Features
A.3.1 Scaling Techniques and Chain-of-Thought
Enhance evaluation with scaling methods and reasoning:
# Self-consistency evaluation
results = evaluator.run(
model="huggingface",
dataset="haerae_bench",
scaling_method=’self_consistency’,
scaling_params={"num_samples": 5}
# CoT with specific trigger phrase
results = evaluator.run(
model="huggingface",
dataset="haerae_bench",
model_params={
"cot": True,
"cot_trigger": "Let’s think step by step."
A.3.2 Model Comparison and Korean Language Support
Compare multiple models and handle language-specific evaluation:
# Define models to compare
models = [
{"model": "huggingface", "model_params": {"model_name_or_path": "Qwen/Qwen2.5-7B-Instruct"}, "name": "Qwen-7B"},
{"model": "huggingface", "model_params": {"model_name_or_path": "kakaocorp/kanana-nano-2.1b-instruct"}, "name": "Kanana-2.1B"}
# Run comparison
comparison_results = evaluator.compare(
models=models,
dataset="haerae_bench",
evaluation_method="string_match"
# Korean language penalizer
results = evaluator.run(
model="huggingface",
dataset="k2_eval",
evaluation_method="string_match",
evaluator_params={
"language_penalizer": True,
"target_language": "ko"
These examples demonstrate HRET’s capabilities for Korean language model evaluation beyond the basic functionality presented in the main paper.
These examples demonstrate HRET’s capabilities for Korean language model evaluation beyond the basic functionality presented in the main paper. Upon publication, comprehensive tutorials and documentation in both Korean and English will be made publicly available in the project repository, including additional examples, best practices, and implementation guidelines.
Appendix B Author Contributions
Hanwool Lee led the project, system design, core development, and quality assurance.
Sooyong Kim contributed extensively to quality assurance, dataset implementation, and scaling methods.
Dasol Choi contributed to paper organization and writing, as well as dataset and evaluation method implementation.
Sangwon Baek worked on backend implementation, tutorial development, and quality assurance.
Seunghyeok Hong contributed to tutorial organization and writing, and implemented several datasets.
Ilgyun Jung developed the LLM-as-a-Judge module and implemented multiple datasets.
Inseon Hwang contributed to dataset implementation and CI/CD pipeline integration.
Naeun Lee worked on backend implementation.
Guijin Son organized the team, contributed to dataset implementation and paper writing, and led the planning and specification of LLM evaluation features.
All authors participated in discussions, testing, and refining the HRET framework.

### v3

@@ 5 Conclusion
We introduce HRET, a unified evaluation framework that addresses critical reproducibility challenges in Korean LLM evaluation. Through its registry-based architecture and standardized protocols, HRET significantly improves evaluation consistency while supporting diverse assessment strategies. HRET integrates major Korean benchmarks with multiple inference backends and evaluation methods, enabling researchers to select task-appropriate strategies while maintaining reproducibility. The framework’s extensible design and community-driven development ensure it evolves with the Korean LLM landscape, providing a sustainable foundation for reproducible assessment.
@@
While HRET provides a standardized evaluation framework for Korean LLMs, several limitations remain. Despite integrating major Korean benchmarks, coverage in specialized domains Hwang et al. (2025) remains limited. The language consistency enforcement may struggle with mixed-language outputs or code-switching common in technical contexts, and computational requirements for test-time scaling techniques can be substantial, potentially limiting accessibility for resource-constrained environments. Current evaluation metrics, particularly string matching, may not fully capture Korean’s rich morphological variations and honorific systems, and the framework’s effectiveness depends on the quality of integrated benchmarks. We are addressing these limitations through ongoing development and welcome community contributions to enhance HRET’s capabilities.
Appendix A Supported Benchmarks
HRET provides unified evaluation support for major Korean language benchmarks spanning diverse domains and task types. The framework’s registry-based architecture enables seamless integration of these benchmarks with consistent preprocessing, evaluation protocols, and result formatting. Table 7 provides detailed descriptions of currently supported benchmarks, organized by evaluation focus.
The benchmark collection covers three primary categories: General benchmarks assess broad Korean language understanding and cultural knowledge, Dialogue benchmarks evaluate conversational capabilities in Korean contexts, and Specialized benchmarks target domain-specific competencies including mathematical reasoning and professional knowledge. This comprehensive coverage enables researchers to conduct thorough evaluations across the spectrum of Korean language model capabilities.
Metric HRET lm-eval-harness
HAE-RAE Bench
Execution Time (sec) 114.93 142.44
Max RAM (MB) 80.96 (<1%) 54.26 (<1%)
Max CPU (%) 525.84 100.0
KMMLU
Execution Time (sec) 303.93 221.04
Max RAM (MB) 104.81 (<1%) 78.76 (<1%)
Max CPU (%) 458.90 199.0
Table 8: Performance comparison on AMD EPYC 74F3 64-Core + NVIDIA H100 with 193GB RAM. Values averaged across 3 runs.
Appendix B Performance Comparison with Existing Tools
Table 8 compares HRET’s computational performance with lm-evaluation-harness using Qwen2.5-3B on identical hardware configurations.
HRET shows mixed performance: faster on HAE-RAE Bench but slower on KMMLU. Higher RAM usage stems from loading datasets as editable objects to enable real-time filtering and experimental modifications through our system. Higher CPU utilization (up to 525%) reflects multi-threading optimization. Both frameworks use <1% of total system memory, indicating efficient resource utilization despite HRET’s flexibility-oriented design.
Appendix C Detailed Reproducibility Results
To complement the summary presented in Section 4.1, we provide the full hardware-wise evaluation results across multiple configurations. This table reports evaluation scores for three representative models using the HuggingFace backend under diverse CPU and GPU settings. These include server-grade (A100, H100) and workstation-grade (RTX A6000, RTX 3090) environments, reflecting real-world variance across institutions.
Model Hardware Setup HAE-RAE CLIcK Batch Size
Qwen2.5-3B AMD EPYC 74F3 + A100×2 0.2847 0.3388 64
AMD EPYC 9354 + RTX A6000 0.2847 0.3388 2
Intel i9-13900K + RTX A6000 0.2847 0.3388 2
Intel Xeon + A100-40GB 0.2863 0.3378 64
Qwen3-8B AMD EPYC 74F3 + A100×2 0.2645 0.3083 64
AMD EPYC 9354 + RTX A6000 0.2645 0.3083 1–2
Intel i9-13900K + RTX A6000 0.2645 0.3083 2
Intel Xeon + A100-40GB 0.2673 0.3083 64
Llama-3.2-3B AMD EPYC 74F3 + A100×2 0.2436 0.2922 64
AMD EPYC 7452 + RTX 3090 0.2437 0.2922 2
Intel i9-13900K + RTX A6000 0.2445 0.2922 2
Intel Xeon + A100-40GB 0.2447 0.2907 64
Table 9: Reproducibility validation across different hardware configurations using HuggingFace backend.
Appendix D Usage Examples and Tutorials
@@
D.1 Advanced Backend Usage
@@
D.1.1 LiteLLM and OpenAI-Compatible APIs
@@
D.2 Evaluation Methods
D.2.1 String Match and Log Probability Evaluation
@@
These examples demonstrate HRET’s capabilities for Korean language model evaluation beyond the basic functionality presented in the main paper. The project repository provides comprehensive tutorials and documentation in both Korean and English, including additional examples, best practices, and implementation guidelines.
Appendix E Real-World Usage
Beyond academic evaluation frameworks, HRET has been adopted in production environments across the Korean NLP industry. Authors’ affiliated organizations, including Shinhan Securities and OneLineAI, have integrated HRET into their model validation and post-training evaluation workflows, demonstrating the framework’s practical utility in real-world deployment scenarios. Additionally, the Haerae Reasoning Challenge (HRC) officially employs HRET to score all participant submissions under standardized protocols, ensuring fairness and transparency in competitive evaluations. This production adoption validates HRET’s reliability and effectiveness beyond controlled research settings, confirming its value for industry practitioners requiring robust evaluation infrastructure.
Ethics Statement
latexHRET aims to improve reproducibility and fairness in Korean LLM evaluation. However, we acknowledge several ethical considerations: (1) standardized evaluation frameworks may inadvertently favor certain model architectures or training approaches; (2) benchmark selection reflects our research perspectives and may not capture all aspects of Korean language competency; (3) language consistency enforcement, while promoting Korean language assessment, may disadvantage multilingual models in certain contexts. We encourage the community to contribute diverse benchmarks and evaluation methods to ensure broad representation of Korean language capabilities. All integrated datasets respect original licensing and guidelines.